# LOLA: LLM-Assisted Online Learning Algorithm for Content Experiments

### v1

@@ Abstract
In the rapidly evolving digital content landscape, media firms and news publishers require automated and efficient methods to enhance user engagement. This paper introduces the LLM-Assisted Online Learning Algorithm (LOLA), a novel framework that integrates Large Language Models (LLMs) with adaptive experimentation to optimize content delivery. Leveraging a large-scale dataset from Upworthy, which includes 17,681 headline A/B tests aimed at evaluating the performance of various headlines associated with the same article content, we first investigate three broad pure-LLM approaches: prompt-based methods, embedding-based classification models, and fine-tuned open-source LLMs. Our findings indicate that prompt-based approaches perform poorly, achieving no more than 65% accuracy in identifying the catchier headline among two options. In contrast, OpenAI-embedding-based classification models and fine-tuned Llama-3-8b models achieve comparable accuracy, around 82-84%, though still falling short of the performance of experimentation with sufficient traffic. We then introduce LOLA, which combines the best pure-LLM approach with the Upper Confidence Bound algorithm to adaptively allocate traffic and maximize clicks. Our numerical experiments on Upworthy data show that LOLA outperforms the standard A/B testing method (the current status quo at Upworthy), pure bandit algorithms, and pure-LLM approaches, particularly in scenarios with limited experimental traffic or numerous arms. Our approach is both scalable and broadly applicable to content experiments across a variety of digital settings where firms seek to optimize user engagement, including digital advertising and social media recommendations.
Keywords: LLMs, Content Experiments, Bandits, News
@@ 1 Introduction
Digital content consumption has seen unprecedented growth, leading to a proliferation of content across various platforms. In today’s real-time digital environment, media firms and news publishers need automated and efficient methods to determine which content generates high user engagement and platform growth. This includes identifying the most appealing articles, the most attractive headlines, and the catchiest cover images (NYTHeadline2017; NYTimes2019). Traditionally, media firms and publishers have relied on experimentation-based approaches to address this problem. Broadly speaking, there are two general types of experimentation styles—(1) Standard A/B tests and (2) Online learning algorithms or bandits. [^1] We use the terms adaptive experimentation, online learning algorithm, and bandits interchangeably throughout the paper. We now discuss each briefly.
@@
A/B testing is the most straightforward experimentation method, where a firm allocates a fixed portion of traffic to different treatment arms, assesses the results, and then goes with the best-performing arm. This is also known as the Explore and Commit strategy, or E&C, where the firm explores for a fixed period using A/B tests and then commits to one treatment based on the results of those tests. This approach is widely used in the publishing industry. For instance, The New York Times recently built a centralized internal A/B test platform, ABRA (A/B Reporting and Allocation architecture), allowing different teams to experiment with their content; see (NYTimes2024) for more details. Another example is Upworthy, which has implemented extensive A/B testing to refine its headlines (matias2021upworthy). Upworthy editors create several headlines for the same article and randomly assign readers to see different headlines with equal probability for a fixed amount of traffic, and then show the best-performing headline to all the users. The main advantage of this approach is that it is trustworthy—with a sufficiently large amount of traffic allocated to the A/B test, the resulting inference is unbiased and accurate. However, a major drawback of this approach is the wastage of traffic—by assigning equal traffic to all the treatment arms during the exploration phase (including the poorly performing ones), the firm incurs high regret or low profits. This is especially problematic in the news industry since news articles tend to have short lifetimes and become stale quickly (typically within a day or two). Therefore, if a firm wastes a lot of traffic to learn the best article/headline, the article itself might become irrelevant by the end of the A/B test.
The second approach, online learning algorithms, or bandits, are able to address some of these problems. These algorithms dynamically adjust traffic allocation to articles/headlines based on their features, past performance, and associated uncertainty. This method provides a more efficient way to optimize content delivery by continually learning and adapting. For instance, both Yahoo News and The New York Times have successfully implemented contextual bandit algorithms to identify content that maximizes user engagement. Yahoo was one of the earliest platforms to use the linear contextual bandit for news article recommendation (li2010contextual). More recently, The New York Times has also adopted a similar algorithm to power its article recommendation system (NYTimes2019).
@@
Online learning algorithms typically outperform A/B tests in terms of overall regret or reward. These algorithms reduce the extent of exploration over time by moving traffic away from poorly performing treatments dynamically. As a result, they tend to switch to the best-performing arms quickly and incur lower regret compared to standard E&C strategies. This has been established both theoretically (lattimore2020bandit) as well empirically. For example, (li2010contextual) demonstrate that the linear contextual bandit algorithm improves the total clicks on news articles over the random policy (i.e., A/B testing with equal probability for arms) at Yahoo News using off-policy evaluation on field data. [^2] Similarly, in the context of digital advertising, (schwartz2017customer) use field experiments to demonstrate that adaptive algorithms outperform the standard balanced A/B test.
@@
Nevertheless, these online algorithms have notable drawbacks. In particular, in standard content experiments, these algorithms typically suffer from a cold-start problem. That is, firms typically start these adaptive experiments under the assumption that all have an equal probability of being the best-performing one. This can lead to a situation where more-than-necessary traffic is allocated to sub-optimal arms and higher regret as less appealing content is shown to users during the exploration phase. Additionally, due to their probabilistic nature, these algorithms can inadvertently converge on a suboptimal arm, especially when the experimental traffic is small, increasing the probability of identifying a suboptimal arm as the best one. These issues arise because online algorithms are designed for large traffic regimes, where the experimenter does not care too much about initialization since it has little impact on the asymptotic regret. However, as discussed earlier, in the news industry, there is limited traffic available for experimentation, given that news content becomes stale quickly. Therefore, the negative effects of the “equally effective” initialization do not diminish sufficiently during the time horizon of interest, leading to higher regret.
@@
Given these issues with content experimentation, the main question that this paper asks and answers is whether we can leverage the recently developed Large Language Models (LLMs) to enhance current content experimentation practices and simplify the process of identifying appealing content. LLMs, trained on extensive human-generated data, have demonstrated the ability to mimic human preferences and behavior in various contexts. For example, (li2024frontiers) investigate whether LLM-generated survey responses produce perceptual maps similar to those generated by human survey responses in car branding analysis. Their study shows a high agreement rate of over 75%. More specifically, in the context of news articles, (yoganarasimhan2024feeds) show that LLMs are effective at understanding the polarization of news content. Hence, a relevant question is whether news publishers can directly use LLMs to predict the appeal of different content, potentially replacing traditional experimentation methods. This approach could significantly reduce the costs associated with experimentation while simplifying the content delivery processes.
@@
To address this question, we leverage a large-scale dataset on content experiments by the publishing firm Upworthy conducted in the 2013–2015 time period. The dataset consists of 17,618 headline tests spanning 77,245 headlines, over 277 million impressions, and over 3.7 million clicks. The goal of each A/B test was to identify the best headline among a set of candidate headlines (for a given article). This dataset is ideal for our study since it was collected from A/B tests where each headline received a relatively large number of impressions, allowing for accurate estimates of each headline’s performance. Such a high-quality dataset provides a trustworthy ground for comparing different approaches. Additionally, since the dataset was sourced from the real media industry, our methods can reflect real-world scenarios and offer practical guidance to both researchers and practitioners.
@@
In the first part of the paper, we use this dataset to examine the extent to which pure-LLM-based approaches can predict headline attractiveness accurately. For this exercise, we focus on pairs of headlines (within a given A/B test) whose CTRs are significantly different (based on the observed data). This gives us a total of 39,158 pairs of headlines and allows us to recast the prediction task as a binary classification task, where a random guess gives 50% accuracy, and the perfect classifier reaches 100% accuracy. Thus, this serves as a preliminary testbed to evaluate whether pure-LLM-based approaches can predict which of the two headlines is catchier (or clickable).
@@
We consider the three most widely used LLM-based approaches for this analysis—(1) Prompt-based approaches, (2) LLM-based text-embedding approaches, and (3) Fine-tuning approaches. We discuss each approach and its performance next. First, for prompt-based methods, we consider two approaches: (a) Zero-shot prompting and (b) In-context learning. Zero-shot prompting is a technique in which an LLM generates responses without being explicitly trained on specific examples. This is the most common way in which most users interact with the LLM. In-context learning involves providing the LLM with several demonstrations of similar tasks before generating responses for a specific task (min2022rethinking; xie2021explanation). We find that GPT-3.5 performs as poorly as a random guess for both prompt-based approaches, with accuracy hovering around 50%. GPT-4 performs marginally better—it achieves an accuracy of 57.26% for zero-shot prompting and 60-65% for in-context learning. Overall, this suggests that prompt-based approaches cannot aid/replace content experiments in any meaningful way.
@@
Next, we turn to the second approach: text embeddings. Specifically, we transform headlines into embedding vectors using the most recent embedding model from OpenAI. We then concatenate two embeddings as inputs, along with the winner with higher CTR as the output, to train binary classification models, including logistic regression and multilayer perceptron (a type of neural network). We find that OpenAI’s embeddings combined with simple logistic regression can achieve around 82% accuracy, while more complex neural networks do not improve performance over the simple logistic model. Finally, we fine-tune a state-of-the-art open-source LLM — Llama-3 with 8 billion parameters, using low-rank adaptation (LoRA) (hu2021lora). This approach performs the best, with 83% accuracy, which is a small improvement over the embedding-based approach. [^3] On the other hand, it also marginally increases training costs and engineering complexity. Overall, we find that prompt-based approaches, which have been the standard method in recent marketing literature for using LLMs so far, perform poorly. In contrast, models based on LLM embeddings or fine-tuned open-source LLMs offer significantly more accurate predictions. The primary advantage of LLM-based methods is that predictions can be made without wasting any traffic. Nonetheless, even the best LLM-based models do not match the accuracy of experimentation-based approaches (with sufficient traffic).
@@
Therefore, in the second part of the paper, we propose and evaluate a novel framework for content experiments that combines the strengths of LLM-based approaches with the advantages of online learning algorithms, termed LOLA (LLM-Assisted Online Learning Algorithm). Our key insight is to use predictions from fine-tuned LLM models as priors in online algorithms to optimize experimentation. The approach involves two steps. First, we train an LLM-based CTR prediction model, which can be an LLM-embedding-based model or a fine-tuned open-source LLM model, or indeed any other LLM-based approach. The second step integrates these predictions into online learning algorithms. LOLA is a general framework, and depending on the specific goals of the experimentation, it can be adapted for different prediction models in the first step and different online algorithms in the second step. For example, if the goal is to minimize regret (i.e., maximize total reward/clicks), we propose a modified Upper Confidence Bound algorithm called LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs), detailed in Section 4.1. The idea of LLM-2UCBs is intuitive: we can view LLM-based CTR predictions as auxiliary samples before the start of the online algorithm (e.g., a 1% CTR can be viewed as 100 impressions with 1 click or 1,000 impressions with 10 clicks). After fine-tuning this auxiliary impression size hyperparameter, we incorporate these imaginary samples into the standard UCB algorithm to get the reward estimator and shrink the upper confidence bound.
@@
LOLA combines the advantages of both experimentation and LLMs, making it accurate and efficient. By using LLMs, we reduce the waste of impressions on poor-performing headlines, especially in the initial stage when we face a cold-start problem and lack good prior information about the headlines’ performance. On the other hand, through experimentation, we can correct the errors in LLMs. Even though LLMs achieve around 82% accuracy, there are still 18% of incorrect cases. This hybrid approach ensures that while LLMs provide reasonable initial predictions, ongoing experimentation refines and improves overall performance, making LOLA an effective solution for content optimization.
@@
We present a comprehensive evaluation of LOLA and compare its performance with three natural benchmarks using the Upworthy dataset: (1) Explore and Commit, (2) pure online learning algorithm, and (3) pure LLM-based model. The first benchmark, E&C, was the status quo at Upworthy during the time of data collection, wherein the editors first ran an A/B test on a set of headlines for a fixed set of impressions, and then displayed the winner for the rest of the traffic. For the pure online learning algorithm, we use the standard Upper Confidence Bound (UCB) algorithm, which is initialized with uniform CTRs. For the pure LLM-based approach we use the linear CTR prediction model in conjunction with OpenAI embeddings. We compute CTRs for all headlines and select the one with the highest CTR for all impressions, bypassing the experimentation phase. [^4] We chose the embedding-based approach for the evaluation instead of the fine-tuned model, given its ease of use and almost comparable performance to the fine-tuned open-source LLM. In LOLA, we use the same CTR prediction model as in the pure LLM approach and employ LLM-2UCBs to maximize accumulated clicks. Our numerical comparison primarily considers the regret minimization framework, as minimizing regret in Upworthy’s context equates to maximizing total clicks over all impressions, including those for experimentation. This metric is pertinent to business profitability since Upworthy, as a free online publisher without subscription fees, relies heavily on advertising revenue. Increased user clicks enable Upworthy to display more advertisements, thereby increasing revenue.
@@
In our empirical evaluation, we find that LOLA outperforms all the baseline approaches. We find that for the earlier impressions, both LOLA and the pure LLM-based approach beat the pure experimentation-based approaches (i.e., the standard A/B test and UCB algorithm) due to relatively accurate LLM-based CTR predictions. However, as the number of impressions grows, the performance of the pure LLM-based approach declines rapidly, as it does not learn anything from past impressions and, as such, cannot improve its performance. Thus, its accuracy falls short compared to other experimentation-based methods. Thus, as discussed earlier, our experiments demonstrate that LOLA is able to leverage the power of LLMs early in the horizon and then build on it to take advantage of experimentation over time, thereby bringing together the strengths of both methods.
Finally, we observe that the performance gap between LOLA and the standard UCB narrows, given that both methods are asymptotically optimal and the benefit provided by LLM diminishes over time. Thus, we expect that LOLA will be especially valuable in settings where the traffic available for experimentation is limited and/or the number of arms is large. Nevertheless, LOLA continues to outperform standard UCB, even after very long horizons.
@@
In summary, our paper makes two key contributions to the literature. First, from a methodological perspective, we develop a novel framework, LOLA, to leverage LLMs for improving content experimentation in digital settings. To the best of our knowledge, this is the first paper that proposes combining LLMs with adaptive experimentation techniques. We demonstrate the value of our approach using a large-scale dataset from Upworthy, showcasing significant improvements in total clicks and profitability. Second, from a managerial perspective, LOLA can be used in various settings where firms need to decide which content to show, such as news articles, headlines, or even advertisements and website content. This method is based on open-source LLM models, making it cost-effective and easily deployable across different domains once fine-tuned on relevant datasets. Compared to the substantial benefits gained from improved content optimization, the cost of fine-tuning these models is negligible. Because this fine-tuning process on the open-sourced LLMs can be efficiently executed even on entry-level GPUs, making it accessible for a wide range of applications. This accessibility ensures that even firms with limited budgets can implement advanced LLM-based techniques to enhance their content strategy. More broadly, while we focused mainly on the news/publishing industry, our framework is general and can easily extend to other settings where firms need to evaluate which type of content maximizes user engagement, including online/digital advertising and social media platforms.

### v2

@@ Abstract
In the rapidly evolving digital content landscape, media firms and news publishers require automated and efficient methods to enhance user engagement. This paper introduces the LLM-Assisted Online Learning Algorithm (LOLA), a novel framework that integrates Large Language Models (LLMs) with adaptive experimentation to optimize content delivery. Leveraging a large-scale dataset from Upworthy, which includes 17,681 headline A/B tests, we first investigate three pure-LLM approaches: prompt-based methods, embedding-based classification models, and fine-tuned open-source LLMs. We find that prompt-based approaches perform poorly, achieving no more than 65% accuracy in identifying the catchier headline. In contrast, both OpenAI-embedding-based classification models and fine-tuned Llama-3 with 8 billion parameters achieve an accuracy of around 82-84%. We then introduce LOLA, which combines the best pure-LLM approach with the Upper Confidence Bound algorithm to allocate traffic and maximize clicks adaptively. Our numerical experiments on Upworthy data show that LOLA outperforms the standard A/B test method (the current status quo at Upworthy), pure bandit algorithms, and pure-LLM approaches, particularly in scenarios with limited experimental traffic. Our approach is scalable and applicable to content experiments across various settings where firms seek to optimize user engagement, including digital advertising and social media recommendations.
Keywords: LLMs, Content Experiments, Bandits, News, Digital marketing.
@@ 1 Introduction
Digital content consumption has seen unprecedented growth, leading to a proliferation of content across various platforms. In today’s real-time digital environment, media firms and news publishers need automated and efficient methods to determine which content generates high user engagement and platform growth. This includes identifying the most appealing articles, the most attractive headlines, and the catchiest cover images (NYTHeadline2017; NYTimes2019). Traditionally, media firms and publishers have relied on experimentation-based approaches to address this problem. Broadly speaking, there are two types of experimentation styles—(1) Standard A/B tests and (2) Online learning algorithms or bandits. [^1] We use the terms adaptive experimentation, online learning algorithm, and bandits interchangeably throughout the paper.
@@
A/B test is the most straightforward experimentation method, where a firm allocates a fixed portion of traffic to different treatment arms, assesses the results, and then goes with the best-performing arm. This is also known as the Explore and Commit strategy, or E&C, where the firm explores for a fixed period using A/B tests and then commits to one treatment based on the results of those tests. This approach is widely used in the publishing industry. For instance, The New York Times recently built a centralized internal A/B test platform, ABRA (A/B Reporting and Allocation architecture), allowing different teams to experiment with their content; see (NYTimes2024) for more details. Another example is Upworthy, which has implemented extensive A/B tests to choose headlines for news articles (matias2021upworthy). The main advantage of this approach is that it is trustworthy—with a sufficiently large amount of traffic allocated to the A/B test, the resulting inference is unbiased and accurate. However, a major drawback of this approach is the wastage of traffic—by assigning equal traffic to all the treatment arms during the exploration phase (including the poorly performing ones), the firm incurs high regret or low profits. This is especially problematic in the news industry since news articles tend to have short lifetimes and become stale quickly (typically within a day or two). Therefore, if a firm wastes a lot of traffic to learn the best article/headline, the article itself might become irrelevant by the end of the A/B test.
@@
The second approach, online learning algorithms, or bandits, are able to address some of these problems. These algorithms dynamically adjust traffic allocation to articles/headlines based on their features, past performance, and associated uncertainty. This method provides a more efficient way to optimize content delivery by continually learning and adapting. For instance, both Yahoo News and The New York Times have successfully implemented contextual bandit algorithms to identify content that maximizes user engagement (li2010contextual; NYTimes2019). Online learning algorithms typically outperform A/B tests in terms of overall regret or reward. These algorithms reduce the extent of exploration over time by moving traffic away from poorly performing treatments dynamically. As a result, they tend to switch to the best-performing arms quickly and incur lower regret compared to standard E&C strategies. This has been established both theoretically (lattimore2020bandit) as well empirically (li2010contextual). [^2] Similarly, in the context of digital advertising, (schwartz2017customer) use field experiments to demonstrate that adaptive algorithms outperform the standard balanced A/B test.
@@
Nevertheless, online algorithms also have notable drawbacks. In standard content experiments, these algorithms suffer from a cold-start problem. That is, firms typically start these adaptive experiments assuming that all arms are equally likely to be the best-performing ones. This can lead to a situation where more-than-necessary traffic is allocated to sub-optimal arms, leading to higher regret. Additionally, due to their probabilistic nature, these algorithms can inadvertently converge on a sub-optimal arm, especially when the experimental traffic is small. These issues arise because online algorithms are designed for large traffic regimes, where the initialization of arms has little impact on the asymptotic regret. However, as discussed earlier, there is limited traffic available for experimentation in the media industry since content becomes stale quickly. Therefore, the negative effects of the “equally effective” initialization do not diminish sufficiently during the time horizon of interest, leading to higher regret.
@@
Given these issues with content experimentation, the main question that this paper asks and answers is whether we can leverage the recently developed Large Language Models (LLMs) to enhance current content experimentation practices and simplify the process of identifying appealing content. LLMs, trained on extensive human-generated data, have demonstrated the ability to mimic human preferences and behavior in a variety of consumer research tasks (li2024frontiers; brand2023using). Specifically, in the context of news articles, (yoganarasimhan2024feeds) show that LLMs can correctly predict the polarization of news content. Hence, a relevant question is whether firms can directly use LLMs to predict the appeal of different content, potentially replacing traditional experimentation. This approach could significantly reduce the costs associated with experimentation while simplifying the content delivery processes.
@@
To address this question, we leverage a large dataset of content experiments by the publishing firm Upworthy conducted in the 2013–2015 timeframe. The dataset consists of 17,618 headline tests spanning 77,245 headlines, 277 million impressions, and 3.7 million clicks. The goal of each A/B test was to identify the best headline among a set of candidate headlines (for a given article). This dataset is an excellent test-bed for our study for a few reasons – (1) it is based on an extremely large number of A/B tests over a high volume of impressions, (2) each A/B test received a relatively large number of impressions, allowing for accurate estimates of each headline’s performance, and (3) it is sourced from a real media firm and represents the actions of a very large number of readers over a sufficiently long period to offer meaningful insights.
@@
In the first part of the paper, we use this dataset to examine the extent to which pure-LLM-based approaches can accurately predict headline attractiveness. For this exercise, we focus on pairs of headlines (within a given A/B test) whose CTRs are significantly different (based on the observed data). This gives us a total of 39,158 pairs of headlines and allows us to recast the prediction task as a binary classification task, where a random guess gives 50% accuracy, and the perfect classifier reaches 100% accuracy. Thus, this serves as a preliminary testbed to evaluate whether pure-LLM-based approaches can predict which of the two headlines is catchier (or clickable).
@@
We consider the three most widely used LLM-based approaches for this analysis—(1) Prompt-based approaches, (2) LLM text-embedding approaches, and (3) Fine-tuning approaches. For prompt-based methods, we consider two approaches: (a) Zero-shot prompting and (b) In-context learning. Zero-shot prompting is a technique in which an LLM generates responses without being explicitly trained on specific examples. This is the most common way in which most users interact with the LLM. In-context learning involves providing the LLM with a few demonstrations of similar tasks before generating responses for a specific task (min2022rethinking; xie2021explanation). We find that GPT-3.5 performs as poorly as a random guess for both prompt-based approaches, with accuracy hovering around 50%. GPT-4 performs marginally better—it achieves an accuracy of 57.26% for zero-shot prompting and 60-65% for in-context learning. Overall, this suggests that prompt-based approaches cannot aid/replace content experiments in any meaningful way.
@@
Next, we turn to the second approach: text embeddings. Specifically, we transform headlines into embedding vectors using the most recent embedding model from OpenAI. We then concatenate two embeddings as inputs, along with the winner with higher CTR as the output, to train binary classification models, including logistic regression and multilayer perceptron (a type of neural network). We find that OpenAI’s embeddings combined with simple logistic regression can achieve around 82.50% accuracy, while more complex neural networks do not improve performance over the simple logistic model. Finally, we fine-tune a state-of-the-art open-source LLM — Llama-3 with 8 billion parameters, using low-rank adaptation (LoRA) (hu2021lora). This approach performs the best, with 83.40% accuracy, which is a small improvement over the embedding-based approach. However, even the best-fine-tuned LLMs are unable to perfectly generate/predict which types of content are most appealing to users. As such, none of the LLM-based approaches can match the accuracy of experimentation-based approaches.
@@
Therefore, in the second part of the paper, we propose and evaluate a novel framework for content experiments that combines the strengths of LLM-based approaches with the advantages of online learning algorithms, termed LOLA (LLM-Assisted Online Learning Algorithm). Our key insight is to use predictions from an LLM model as priors in online algorithms to optimize experimentation. The approach involves two steps. First, we train an LLM-based CTR prediction model using LoRA fine-tuning. The second step integrates these predictions into online learning algorithms. LOLA is a general framework, and depending on the specific goals of the experimentation, it can be adapted for different prediction models in the first step and different online algorithms in the second step. For example, if the goal is to minimize regret (i.e., maximize total reward/clicks), we propose a modified Upper Confidence Bound algorithm called LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs) that is detailed in §$\S$4.2. The idea of LLM-2UCBs is intuitive: we can view LLM-based CTR predictions as auxiliary samples before the start of the online algorithm (e.g., a 1% CTR can be viewed as 100 impressions with 1 click or 1,000 impressions with 10 clicks). After fine-tuning this auxiliary impression size hyperparameter, we incorporate these auxiliary samples into the standard UCB algorithm (auer2002finite) to obtain the reward estimator and shrink the upper confidence bound.
@@
LOLA combines the advantages of both experimentation and LLMs, making it accurate and efficient. By integrating an LLM-based model to predict CTRs before the online deployment phase, we avoid wasting impressions on poorly performing headlines, especially in the initial stage when we face a cold-start problem. On the other hand, through experimentation, we can correct the errors in LLM predictions. This hybrid approach ensures that while LLMs provide reasonable initial predictions, ongoing experimentation refines and improves overall performance, making LOLA an effective solution for optimizing content delivery.
@@
We present a comprehensive evaluation of LOLA and compare its performance with three natural benchmarks using the Upworthy dataset: (1) Explore and Commit, (2) a pure online learning algorithm, and (3) a pure LLM-based model. The first benchmark, E&C, was the status quo at Upworthy during the time of data collection, wherein the editors first ran an A/B test on a set of headlines for a fixed set of impressions and then displayed the winner for the rest of the traffic. For the pure online learning algorithm, we use the standard Upper Confidence Bound (UCB) algorithm, which is initialized with uniform CTRs. For the pure LLM-based approach, we use the LoRA fine-tuned Llama-3 to predict the CTRs of headlines and select the one with the highest CTR for all impressions, bypassing the experimentation phase. In LOLA, we use the same CTR prediction model as that used in the pure LLM approach and then employ LLM-2UCBs to maximize accumulated clicks. We find that LOLA outperforms all the baseline approaches, though the next-best algorithm varies based on the length of the time horizon (keeping the number of headlines fixed).
@@
Specifically, when time horizons are small, both LOLA and the pure LLM-based approach beat the pure experimentation-based approaches (i.e., E&C and the UCB algorithm) due to relatively accurate LLM-based CTR predictions, and lack of sufficient time for the experimentation to provide meaningful updates. Specifically, both LOLA and the pure LLM outperform E&C by 8-9% and UCB by 6-7% in this scenario. However, as the number of impressions/time horizon grows, the performance of the pure LLM-based approach falls behind other experimentation-based methods because pure LLM-based approaches are static and do not take advantage of experimentation. In this case, LOLA beats pure LLM methods and E&C by 4–5%, and UCB by 2-3%. In sum, our experiments demonstrate that LOLA is able to leverage the power of LLMs early in the horizon and then build on it to take advantage of experimentation over time, thereby bringing together the strengths of both methods.
Finally, while we focus on regret minimization in a stochastic bandit setting given the Upworthy context, LOLA is a general-purpose framework and can easily accommodate different goals and/or settings. We provide variants of the standard LOLA (and results on the empirical performance of these variants when applicable) for a few natural extensions, e.g., for Bayesian bandit settings where the second stage algorithm can be Thompson Sampling instead of UCB, for the best arm identification problem where the goal is to identify the best content/headline rather than regret minimization, and for settings where the firm has access to user/contextual features that can be used to personalize content delivery.
@@
In summary, our paper makes two key contributions to the literature. First, from a methodological perspective, we develop a novel framework, LOLA, to leverage LLMs for improving content experimentation in digital settings. To the best of our knowledge, this is the first paper that proposes combining LLMs with adaptive experimentation techniques. We demonstrate the value of our approach using a large-scale dataset from Upworthy when compared to baseline methods. Second, from a managerial perspective, LOLA can be used in a broad range of settings where firms need to decide which content to show. While we focus on the news/publishing industry given our empirical context, the LOLA framework is general and can easily extend to other settings, including digital advertising, email marketing, and website design. Further, the method is cost-effective, can be built on open-source LLMs (alleviating data privacy concerns), and is easily deployable across different domains once fine-tuned on relevant datasets.